OpenStudy (anonymous):

solve the following: 10^(2x-3) = 0.01

OpenStudy (klimenkov):

If \(a^x=b^x\), then \(a=b\). And use \(0.01=\frac1 {100}=\frac1{10^2}=10^{-2}\).

OpenStudy (klimenkov):

Oops. Really you need is that if \(a^x=a^y\) then \(x=y\).

OpenStudy (anonymous):

Use what @klimenkov said above. \[10^{2x - 3} = 0.01\]\[10^{2x - 3} = \frac{1}{100}\]\[10^{2x - 3} = \frac{1}{10^{2}}\]\[10^{2x - 3} = 10^{-2}\]\[2x - 3 = -2\]Solve for x now.

OpenStudy (klimenkov):

The very first statement is false. Forget it, because if \(2^2=(-2)^2\), but \(2\ne-2\).

OpenStudy (anonymous):

What do you mean? I'm not quite following.

OpenStudy (klimenkov):

I said that mine statement was false.

OpenStudy (anonymous):

Oh ok.

OpenStudy (anonymous):

i'm sorry. which statement was false?

OpenStudy (anonymous):

would the answer be x = 1/2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I am so late...but you are correct :)